OpenStudy (anonymous):

Solve ((x+2)^2+16^2)^(1/2)/[16^2(x+2)]=

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

let me know if you're here and need help. is that the exact question as written?

OpenStudy (anonymous):

√(u^2+a^2)/(a^2*u)= u=x+2 and a=16

OpenStudy (anonymous):

Hey in the denominator, is it \[A ^{2}U\] or \[A ^{2U}\]

OpenStudy (anonymous):

Ive got \[\left(\sqrt{U ^{2}+A ^{2}} \right)/A ^{2}U\]

OpenStudy (anonymous):

[√((x+2)^2+16^2)]/(16^2(x+2))=

OpenStudy (anonymous):

I have a problem calculating this

OpenStudy (anonymous):

same...